Determine the number of hydrogen ions in 56cm^3 of solution of pH 3.
1
Expert's answer
2019-10-14T14:31:16-0400

Solution.

Since we do not specify the name of the acid, then take for a fact that the acid is strong, and the formula for pH is as follows:

"pH = -lg[H^+]"

"[H^+] = 10^{-pH}"

[H+] = 0.001 M

"C = \frac{n}{V}"

"n = C \times V"

"n(H^+) = 0.001 \times 0.056 = 5.6*10^{-5} \ mole"

Answer:

"n(H^+) = 5.6*10^{-5} \ mole"